OpenStudy (just_gymnast_me):

Trevor is tiling his bathroom floor, which has an area that is represented as 120r3 square inches. Each tile has an area of the square root of the quantity 16 times r to the ninth power.. Use the given information to write and simplify an expression for the total number of tiles used.

OpenStudy (just_gymnast_me):

not really sure how to do this..please help! ;/

OpenStudy (just_gymnast_me):

thank you! can you explain how to get the answer a little bit?

OpenStudy (just_gymnast_me):

@bohotness

OpenStudy (just_gymnast_me):

im still confused..

OpenStudy (anonymous):

ANSWER: TILES USED = 30 / [r * sqrt(r)] Why??? TILES USED = TOTAL AREA / AREA per TILE = 120r^3 / sqrt(16r^9) = 120r^3 / [sqrt(16) * sqrt(r^9)] =120r^3 / [4 * sqrt(r^8) *sqrt(r)] = 120r^3 / [ 4 * r^4 * sqrt(4)] =30 / [r * sqrt(r)]